Greg invested a sum of money at 9.5% APR, compounded monthly. He let the sum accumulate interest for 25 years and now has $31,952.82. What was Greg's initial amount that he invested?
Answer #1

This question requires application of basic time value of money function,

FV = PV * (1 + r)n

Now, since the compounding is happening on monthly basis,

r = 9.5%/12 = 0.7917%

n = 25 * 12 = 300 months

FV = $31,952.82

$31,952.82 = PV * (1 + 0.007917)300

$31,952.82 = PV * 10.6509

PV = $3000